OpenStudy (anonymous):

Sequences 3125,625,125,25,5,...

OpenStudy (anonymous):

The sequence is geometric. Well what pattern can you see with numbers? How many times bigger is 125 compared to 25. How many times bigger is 625 compared to 125? Find this relationship then you should be able to find the next term...

OpenStudy (akashdeepdeb):

Do you see any relation? 3125/5 = 625 625/5 = 125 125/5 = 25 25/5 = 5 Getting anything from this? :D

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."